### v1

@@ Abstract
The accelerating loss of biodiversity presents critical challenges for ecological research and conservation strategies. The preservation of biodiversity is paramount for maintaining ecological balance and ensuring the sustainability of ecosystems. However, biodiversity faces numerous threats, including habitat loss, climate change, and the proliferation of invasive species. Addressing these and other ecology-related challenges, both at local and global scales, requires comprehensive monitoring, predictive and conservation planning capabilities. Artificial Intelligence (AI) Foundation Models (FMs) have gained significant momentum in numerous scientific domains by leveraging vast datasets to learn general-purpose representations adaptable to various downstream tasks. This paradigm holds immense promise for biodiversity conservation. In response, we introduce BioAnalyst, the first Foundation Model tailored for biodiversity analysis and conservation planning. BioAnalyst employs a transformer-based architecture, pre-trained on extensive multi-modal datasets encompassing species occurrence records, remote sensing indicators, climate and environmental variables. BioAnalyst is designed for adaptability, allowing for fine-tuning of a range of downstream tasks, such as species distribution modelling, habitat suitability assessments, invasive species detection, and population trend forecasting. We evaluate the model’s performance on two downstream use cases, demonstrating its generalisability compared to existing methods, particularly in data-scarce scenarios for two distinct use-cases, establishing a new accuracy baseline for ecological forecasting. By openly releasing BioAnalyst and its fine-tuning workflows to the scientific community, we aim to foster collaborative efforts in biodiversity modelling and advance AI-driven solutions to pressing ecological challenges.
keywords:
Foundation Model, Deep Learning, Representation Learning, Ecology, Biodiversity
@@ 1 Introduction
Biodiversity, encompassing the variety of all life forms on Earth, is fundamental to the stability and resilience of ecosystems. However, this rich diversity is under unprecedented threat due to numerous factors such as climate change [1], pollution [2, 3], habitat destruction, over-exploitation of natural resources [4], and the introduction of invasive species [5]. These pressures have led to significant declines in species populations and ecosystem degradation, posing critical risks to human well-being by compromising essential ecosystem services, such as clean air, water, and fertile soil [6].
@@
Addressing these challenges requires predictive models to understand ecosystem dynamics and quantify the impacts of interventions. This, in turn, raises the overarching question of how to integrate such insights into decision-support frameworks for biodiversity conservation. Traditional methods often rely on static models, such as species distribution maps [7], which lack real-time updates and fail to capture rapid environmental changes. The fragmented nature of biodiversity data, dispersed across various sources and formats [8], hinders effective data harmonisation and integration. Additionally, ecological systems are inherently complex and less understood compared to engineered systems, making accurate modelling and prediction arduous tasks. Uncertainties and knowledge gaps persist, particularly in identifying and accounting for unknown variables and intricate inter-species interactions [9].
@@
Recent advancements in AI and the development of Foundation Models offer promising avenues to overcome these challenges [10]. FMs, pre-trained on large-scale datasets primarily through self-supervision, have revolutionised fields such as natural language processing [11] and computer vision [12], demonstrating remarkable adaptability across diverse tasks. While geospatial foundation models are increasingly applied in ecological research, biodiversity modelling presents distinct challenges due to its reliance on unique data modalities, such as species occurrence records, trait databases, and fine-scale environmental covariates, for which specialised FMs have yet to be developed. The complexity and heterogeneity of ecological data, including species occurrence records, genetic sequences, remote sensing imagery, climate data, and environmental variables, pose significant challenges for integration and scalability. Moreover, the lack of standardised protocols for data collection and model development further complicates the creation of comprehensive AI tools in this domain [13].
@@
In response to these challenges, we introduce BioAnalyst, a Foundation Model specifically designed for biodiversity analytics, opening new avenues on both local and global scale conservation planning efforts. Our contributions in this work are threefold:
@@
  • Development of the first Multi-modal Biodiversity Foundation Model: We present BioAnalyst as the first large-scale AI model tailored for biodiversity modelling, capable of processing and integrating diverse data types to model complex ecological phenomena.
@@ 2 Related Work
One of the first successful applications of FMs in Earth Sciences involves a geospatial foundation model trained on raw satellite imagery, called Prithvi, which can tackle tasks such as flood mapping, wildfire scar segmentation, multi-temporal crop segmentation, and cloud gap imputation [14]. A follow-up work introduced Prithvi WxC, a larger 2.2 billion-parameter FM that emulates weather and climate phenomena in tasks such as autoregressive rollout forecasting, downscaling, gravity wave parameterisation, and extreme events estimation [15]. On the same theme, Pangu-Weather delivered higher performance in medium-range forecasting, improving numeric weather prediction methods by training a 3D transformer model on 39 years of global data, which injects Earth-specific priors [16].
@@
Focusing on Earth system dynamics and predictability, as well as the more specific and accurate prediction of extreme weather and climate events, ORBIT showcased advanced performance and highlighted the requirement for High Performance Computing (HPC) [17]. Similarly, Aurora can produce operational forecasts for global medium-range weather with unprecedented accuracy and speed-up over classical numerical weather prediction (NWP) models, by combining a flexible 3D Transformer backbone with a distinct encoder-decoder architecture [18]. Following similar approaches, Aardvak Weather features an end-to-end pipeline for data-driven weather prediction focusing on computation and maintenance benefits compared to classic NWP models [19]. In the Earth Observation domain, TerraMind is a large FM pre-trained on nine distinct modalities, highlighting the powerful alignment of token and pixel-level representations. In addition, this work demonstrates both the benefits of early fusion on downstream tasks and the performance gains achieved when learning on modalities generated by the FM [20]. To stimulate the development of FMs for earth monitoring, GEO-bench offers a suite of six classification and six segmentation tasks, suited for model evaluation [21].
@@
In ecology, the number of FMs is relatively small, with a focus on visual, audio, and natural language tasks. More specifically, BioCLIP is an FM classifier for biology for the tree of life, trained on the TREEOFLIFE-10M, namely the abundance and variety of images of plants, animals, and fungi, together with the availability of rich structured biological knowledge [22]. Similarly, Insect-Foundation introduced a 1M dataset with insect imagery and taxonomy, and an FM based on ViT backbone [12] trained on this dataset for classification [23]. For species distribution modelling, NicheFlow demonstrated good predictive performance, mainly in reptile species [24], employing a Variational Autoencoder architecture and using environmental and species distribution variables. Combining audio and textual information NatureLM uses a pretrained encoder and a frozen LLM backbone (Llama 3.1-8b) to produce a text sequence used for bioacoustics tasks and more specific species-classification and detection [25].
@@ 3 Method
BioAnalyst has been designed to utilise the predictive power of the latest AI transformer-based models while being flexible enough to digest multi-modal geospatial input variables. Our work is inspired by the development of large-scale climate and weather models, such as Aurora [18] and Prithvi [15, 14], extending the capabilities of Foundation Models in the domain of biodiversity. More specifically, we are interested in learning about and forecasting biodiversity dynamics at both global and local scales with adequate resolution.
The design choices of BioAnalyst were driven by specific capabilities that it should possess, including multi-modal data representation, spatiotemporal feature preservation, global and local scale operation, underlying physics simulation across multiple scales, and various use-case adaptability. To account for them, BioAnalyst can be thought of as a forecast emulator, i.e., given a state of the Earth’s biodiversity at times $t$ and $t-\delta t$, it predicts the state at $t+\delta t$, where $\delta$ is a discrete time step. Although this might seem very simple, it poses significant challenges for modelling and engineering in complex domains such as ecology and, more specifically, biodiversity, which we have attempted to tackle to the best of our ability given the constrained resources at hand.
@@ 3.1 Foundation Model Architecture
Forecasting is a common task in Earth Sciences, such as weather, climate, and ecology, which is mainly modelled with time-series methods. Related work on weather and climate utilises the latest advancements in computer vision literature, including masked autoencoders [26], which exploit their low memory footprint, masking properties, and the handling of ungridded and sparse observation data.
@@
  • Encoder $\mathcal{E}$: We use Perceiver IO [27, 28], a general-purpose attention architecture. Input variables $\mathbf{X}_{t}$ are first tokenized into $N_{p}=\mathcal{H}/p\times\mathcal{W}/p$ non-overlapping patches of size $p\times p$. Fourier features encode spatial coordinates, which are combined with learned embeddings for variable types, time steps, and atmospheric levels. The resulting features associated with each patch are projected into the model’s embedding dimension $D_{e}$, creating tokens $\mathbf{T}_{t}\in\mathbb{R}^{N_{p}\times D_{e}}$. These are processed by the Perceiver IO encoder $\mathcal{E}$, which maps them to a fixed-size latent array $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$ (where $N_{l}$ is the number of latent tokens) using cross-attention followed by self-attention layers:
@@
  • Backbone $\mathcal{B}$: We use a SwinTransformer [29] as the neural simulation engine. It receives the latent representations from two previous steps $\mathbf{Z}_{t-1},\mathbf{Z}_{t}$ and predicts the next latent state $\mathbf{Z}^{\prime}_{t+1}$ using hierarchical stages with shifted window self-attention:
@@
[…]
The design choices for BioAnalyst prioritise learning informative features in a compact latent representation before the emulation stage. By using Perceiver IO for both encoding and decoding stages, we aim to learn features from the original (raw) input data, thereby avoiding the standard approach of using separate tokenisation pipelines for each modality/variable-type, which can lead to biased tokens that are heavily dependent on the model used to produce them. This unified approach enables the model to capture cross-modal interactions at different data granularity levels, allowing it to differentiate features across various domains, from ground conditions and atmospheric levels to species distributions.
@@ 3.2 Pre-training Data Selection
BioAnalyst is pre-trained on BioCube [30], which compiles and aligns multiple datasets into a fixed spatio-temporal cube. Our main driver is modelling biodiversity ”as-a-whole”, which means we require observations from below the surface, the surface and above. Our analysis is confined to terrestrial (land-based) biodiversity; therefore, datasets describing marine or coastal biota are intentionally excluded from the study. More specifically, we select a subset of the total available features, categorised by modality groups, namely:
  • Atmospheric variables with 13 levels
  • Climate variables
  • Edaphic variables
  • Vegetation variables
  • Species Distribution variables
  • Land, Agriculture and Forest variables
  • RedList variables
  • Miscellaneous variables
These features are combined in a Data HyperCube, grounded on the coordinate reference system WGS84. The HyperCube contains global coordinates with a resolution of 0.25 degrees (grid sampling $\sim 28$ km) from the whole world while our focus is on European biodiversity, leading us to select a slice from it, yielding a Data Batch from [latitude: $\mathcal{H}$, longitude: $\mathcal{W}$ ] = [(32, 72), (-25, 45)] = [160, 280]. The observation time range spans from January 1, 2000, to June 1, 2020, and we sample with a 1-month lead time from this range.
@@
Selecting a Data Sample from the Data Batch yields a composite multi-modal cell of European coordinates, with a specific monthly time-stamp. Each of these Data Points contains a total of 113 observations per location cell. More specifically, we denote the observed data points at a discrete time $t$ by a collection of tensors $\mathbf{X_{t}}$:
$\mathbf{X}_{t}=\sum_{i=1}^{10}\sum_{j=1}^{13}variables_{i}\cdot(levels_{j}),\mathbf{X}_{t}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times 113}$ (4)

### v2

@@ Abstract
Multimodal Foundation Models (FMs) offer a path to learn general-purpose representations from heterogeneous ecological data, easily transferable to downstream tasks. However, practical biodiversity modelling remains fragmented; separate pipelines and models are built for each dataset and objective, which limits reuse across regions and taxa. In response, we present BioAnalyst, to our knowledge the first multimodal Foundation Model tailored to biodiversity analysis and conservation planning in Europe at $0.25^{\circ}$ spatial resolution targeting regional to national-scale applications. BioAnalyst employs a transformer-based architecture, pre-trained on extensive multimodal datasets that align species occurrence records with remote sensing indicators, climate and environmental variables. Post pre-training, the model is adapted via lightweight roll-out fine-tuning to a range of downstream tasks, including joint species distribution modelling, biodiversity dynamics and population trend forecasting. The model is evaluated on two representative downstream use cases: (i) joint species distribution modelling and with 500 vascular plant species (ii) monthly climate linear probing with temperature and precipitation data. Our findings show that BioAnalyst can provide a strong baseline both for biotic and abiotic tasks, acting as a macroecological simulator with a yearly forecasting horizon and monthly resolution, offering the first application of this type of modelling in the biodiversity domain. We have open-sourced the model weights, training and fine-tuning pipelines to advance AI-driven ecological research.
@@ 1 Introduction
Biodiversity, encompassing the variety of all life forms on Earth, is fundamental to the stability and resilience of ecosystems. However, this rich diversity is under unprecedented threat due to numerous factors such as climate change (Gitay2003ClimateCB), pollution (Manisalidis2020EnvironmentalAH; AzevedoSantos2021PlasticPA), habitat destruction, over-exploitation of natural resources (Cordes2016EnvironmentalIO), and the introduction of invasive species (CrystalOrnelas2020TheU). These pressures have led to significant declines in species populations and ecosystem degradation, posing critical risks to human well-being by compromising essential ecosystem services, such as clean air, water, and fertile soil (Cardinale2012BiodiversityLA).
@@
Addressing these challenges requires predictive models to understand ecosystem dynamics and quantify the impacts of interventions. This, in turn, raises the overarching question of how to integrate such insights into decision-support frameworks for biodiversity conservation. Traditional methods often rely on static models, such as species distribution maps (JUNG2023102127), which lack real-time updates and fail to capture rapid environmental changes. The fragmented nature of biodiversity data, dispersed across various sources and formats (WOHNER2022101708), hinders effective data harmonisation and integration. Additionally, ecological systems are inherently complex and less understood compared to engineered systems, making accurate modelling and prediction arduous tasks. Uncertainties and knowledge gaps persist, particularly in identifying and accounting for unknown variables and intricate inter-species interactions (ZHU2022101760).
@@
Recent advancements in AI and the development of Foundation Models offer promising avenues to overcome these challenges (bommasani2021opportunities). FMs, pre-trained on large-scale datasets primarily through self-supervision, have revolutionised fields such as natural language processing (NEURIPS2020_1457c0d6) and computer vision (dosovitskiy2020image), demonstrating remarkable adaptability across diverse tasks. While geospatial foundation models are increasingly applied in ecological research, biodiversity modelling presents distinct challenges due to its reliance on unique data modalities, such as species occurrence records, trait databases, and fine-scale environmental covariates, for which specialised FMs have yet to be developed. The complexity and heterogeneity of ecological data, including species occurrence records, genetic sequences, remote sensing imagery, climate data, and environmental variables, pose significant challenges for integration and scalability. Moreover, the lack of standardised protocols for data collection and model development further complicates the creation of comprehensive AI tools in this domain (TRANTAS2023102357).
@@
In response to these challenges, we introduce BioAnalyst, a Foundation Model specifically designed for biodiversity analytics, opening new avenues on both regional and national scale conservation planning efforts. Our contributions in this work are threefold:
@@
  • Development of the first Multi-modal Biodiversity Foundation Model: To our knowledge the first large-scale AI model tailored for biodiversity modelling, capable of processing and integrating diverse data types to model complex ecological phenomena.
@@ 2 Related Work
One of the first successful applications of FMs in Earth Sciences involves a geospatial foundation model trained on raw satellite imagery, called Prithvi, which can tackle tasks such as flood mapping, wildfire scar segmentation, multi-temporal crop segmentation, and cloud gap imputation (Jakubik2023FoundationMF). A follow-up work introduced Prithvi WxC, a larger 2.2 billion-parameter FM that emulates weather and climate phenomena in tasks such as autoregressive rollout forecasting, downscaling, gravity wave parameterisation, and extreme events estimation (schmude2024prithviwxcfoundationmodel). On the same theme, Pangu-Weather delivered higher performance in medium-range forecasting, improving numeric weather prediction methods by training a 3D transformer model on 39 years of global data, which injects Earth-specific priors (bi2023accurate).
@@
Focusing on Earth system dynamics and predictability, as well as the more specific and accurate prediction of extreme weather and climate events, ORBIT showcased advanced performance and highlighted the requirement for High Performance Computing (HPC) (wang2024orbit). Similarly, Aurora can produce operational forecasts for global medium-range weather with unprecedented accuracy and speed-up over classical numerical weather prediction (NWP) models, by combining a flexible 3D Transformer backbone with a distinct encoder-decoder architecture (bodnar2024aurora). Following similar approaches, Aardvak Weather features an end-to-end pipeline for data-driven weather prediction focusing on computation and maintenance benefits compared to classic NWP models (allen2025end). In the Earth Observation domain, TerraMind is a large FM pre-trained on nine distinct modalities, highlighting the powerful alignment of token and pixel-level representations. In addition, this work demonstrates both the benefits of early fusion on downstream tasks and the performance gains achieved when learning on modalities generated by the FM (jakubik2025terramindlargescalegenerativemultimodality). To stimulate the development of FMs for earth monitoring, GEO-bench offers a suite of six classification and six segmentation tasks, suited for model evaluation (lacoste2024geo).
@@
In ecology, the number of FMs is relatively small, with a focus on visual, audio, and natural language tasks. More specifically, BioCLIP is an FM classifier for biology for the tree of life, trained on the TREEOFLIFE-10M, namely the abundance and variety of images of plants, animals, and fungi, together with the availability of rich structured biological knowledge (stevens2024bioclip). Similarly, Insect-Foundation introduced a 1M dataset with insect imagery and taxonomy, and an FM based on ViT backbone (dosovitskiy2020image) trained on this dataset for classification (nguyen2024insect). For species distribution modelling, NicheFlow demonstrated good predictive performance, mainly in reptile species (Dinnage2024.10.15.618541), employing a Variational Autoencoder architecture and using environmental and species distribution variables. Combining audio and textual information NatureLM uses a pretrained encoder and a frozen LLM backbone (Llama 3.1-8b) to produce a text sequence used for bioacoustics tasks and more specific species-classification and detection (robinson2024naturelm-audio).
@@ 3 Method
BioAnalyst has been designed to utilise the predictive power of the latest AI transformer-based models while being flexible enough to digest multi-modal geospatial input variables. Our work is inspired by the development of large-scale climate and weather models, such as Aurora (bodnar2024aurora) and Prithvi (schmude2024prithviwxcfoundationmodel; Jakubik2023FoundationMF), extending the capabilities of Foundation Models in the domain of biodiversity. More specifically, we are interested in learning about and forecasting biodiversity dynamics at both regional and national scales with adequate resolution.
The design choices of BioAnalyst were driven by specific capabilities that should possess, including multi-modal data representation, spatiotemporal feature preservation, regional and national operation, underlying physics simulation across multiple scales, and various use-case adaptability. We selected 28 km resolution for BioAnalyst, as it is appropriate for regional to national-scale applications such as (i) mapping broad richness and community-composition patterns, (ii) identifying large-scale hotspots and coldspots under different climate scenarios, and (iii) informing high-level prioritization or reporting (e.g. national assessments, EU-wide strategies). To account for them, BioAnalyst can be thought of as a forecast emulator, i.e., given a state of the Earth’s biodiversity at times $t$ and $t-\delta t$, it predicts the state at $t+\delta t$, where $\delta$ is a discrete time step and the final fine-tuned model can predict up to 12 time-steps ahead. Although this might seem very simple, it poses significant challenges for modelling and engineering in complex domains such as ecology and, more specifically, biodiversity, which we have attempted to tackle to the best of our ability given the constrained resources at hand.
@@ 3.1 Foundation Model Architecture
Forecasting is a common task in Earth Sciences, such as weather, climate, and ecology, which is mainly modelled with time-series methods. Related work on weather and climate utilises the latest advancements in computer vision literature, including masked autoencoders (he2022masked), which exploit their low memory footprint, masking properties, and the handling of ungridded and sparse observation data.
@@
  • Encoder $\mathcal{E}$: We use Perceiver IO (jaegle2021perceiver; jaegle2021perceiverio), a general-purpose attention architecture. Input variables $\mathbf{X}_{t}$ are first tokenized into $N_{p}=\mathcal{H}/p\times\mathcal{W}/p$ non-overlapping patches of size $p\times p$. Fourier features encode spatial coordinates, which are combined with learned embeddings for variable types, time steps, and atmospheric levels. The resulting features associated with each patch are projected into the model’s embedding dimension $D_{e}$, creating tokens $\mathbf{T}_{t}\in\mathbb{R}^{N_{p}\times D_{e}}$. These are processed by the Perceiver IO encoder $\mathcal{E}$, which maps them to a fixed-size latent array $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$ (where $N_{l}$ is the number of latent tokens) using cross-attention followed by self-attention layers:
@@
  • Backbone $\mathcal{B}$: We use a SwinTransformer (liu2021swin) as the neural simulation engine. It receives the latent representations from two previous steps $\mathbf{Z}_{t-1},\mathbf{Z}_{t}$ and predicts the next latent state $\mathbf{Z}^{\prime}_{t+1}$ using hierarchical stages with shifted window self-attention:
@@
    $\mathbf{\hat{Y}}_{t+1}=\mathcal{D}(\mathbf{Z}^{\prime}_{t+1},\mathbf{Q})\quad\xrightarrow{\text{Reshape}}\quad\mathbf{\hat{X}}_{t+1}$ (3)
The design choices for BioAnalyst prioritise learning informative features in a compact latent representation before the emulation stage. By using Perceiver IO for both encoding and decoding stages, we aim to learn features from the original (raw) input data, thereby avoiding the standard approach of using separate tokenisation pipelines for each modality/variable-type, which can lead to biased tokens that are heavily dependent on the model used to produce them. This unified approach enables the model to capture cross-modal interactions at different data granularity levels, allowing it to differentiate features across various domains, from ground conditions and atmospheric levels to species distributions. More detailed information on the architecture is available in Appendix A.
@@ 3.2 Pre-training Data Selection
BioAnalyst is pre-trained on BioCube (stasinos2025biocube), which compiles and aligns multiple datasets into a fixed spatio-temporal cube. Our main driver is modelling biodiversity “as-a-whole”, which means we require observations from below the surface, the surface and above. Our analysis is confined to terrestrial (land-based) biodiversity; therefore, datasets describing marine or coastal biota are intentionally excluded from the study. Additionally, we select a subset of the total available features, categorised by variable groups, namely: surface variables, atmospheric variables with 13 levels, climate variables, edaphic variables, vegetation variables, species distribution variables, land, agriculture and forest variables, redlist variable and miscellaneous variables. Our species distribution variables consist of animal species only and not plants.
These features are combined in a Data HyperCube, grounded on the coordinate reference system WGS84. The HyperCube contains global coordinates with a resolution of 0.25 degrees (grid sampling $\sim 28$ km) from the whole world while our focus is on European biodiversity, leading us to select a slice from it, yielding a Data Batch from [latitude: $\mathcal{H}$, longitude: $\mathcal{W}$ ] = [(32, 72), (-25, 45)] = [160, 280]. The observation time range spans from January 1, 2000, to June 1, 2020, and we sample with a 1-month lead time from this range, resulting in total to 233 monthly Data Samples.
@@
Selecting a Data Sample from the Data Batch yields a composite multi-modal cell of European coordinates, with a specific monthly time-stamp. Each of these Data Points for $t\in[0,233]$ contains 124 observations per location cell. The total observations are calculated by summing the 11 variable groups with the number of variables they contain and for the atmospheric variables multiplying the variables with the number of pressure levels and adding them to the sum. More specifically, we denote the observed data points at a discrete time $t$ by a collection of tensors $\mathbf{X}_{t}$:
$\mathbf{X}_{t}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times F},$ (4)
with $\mathcal{H}=160$, $\mathcal{W}=280$, and $F=124$ feature channels per spatial cell. We consider $I=11$ variable groups indexed by $i=1,\dots,I$, where the $i$-th group at time $t$ is represented as
$V_{t}^{(i)}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times L_{i}},$ (5)
with $L_{i}$ the number of levels associated with that group (e.g. 13 pressure levels for atmospheric variables or a single level for all other variable groups). A full Data Sample is obtained by concatenating all variable-group tensors along the feature dimension,
$\mathbf{X}_{t}=\operatorname{concat}_{i=1}^{I}V_{t}^{(i)},$ (6)
so that each spatial cell $(\mathcal{H},\mathcal{W})$ at time $t$ is associated with a feature vector $x_{t,h,w}^{v}\in\mathbb{R}^{124},v\in V$.
During pre-training and roll-out fine-tuning, the input data consist of tuples of two consecutive time-steps. A generic input at time index $t$ is given by
$\bigl(\mathbf{X}_{t},\mathbf{X}_{t+1}\bigr)\in\left(\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times F}\right)^{2},$ (7)